I need help with this problem I am using 2l+2w but the answer I am getting isn't correct. Can someone please help me to understand what I am doing wrong.
The question or problem I am trying to solve is:
The length of a rectangle is 9 centimeters more than it's width. The perimeter is 82 centimeters. Find the demensions of the rectangle.
Please help me with this problem. I am sleepy and getting very upset at this time.

Answer by ankor@dixie-net.com(22740)

The length of a rectangle is 9 centimeters more than it's width.
This says: L = (W+9)
:
The perimeter is 82 centimeters.
you know: 2L + 2W = 82
;
Substitute (W+9) for L, and find W:
2(W+9) + 2W = 82
:
Simplify, divide equation by 2 and you have:
W + 9 + W = 41
2W = 41 - 9
2W = 32
W = 32/2
W = 16 is the width
:
It said the length was 9 cm more: L = 25 cm
:
:
Check the perimeter now
2(25) + 2(16) = 82